### v1

# Attention Via Convolutional Nearest Neighbors

Abstract
The shift from Convolutional Neural Networks to Transformers has reshaped computer vision, yet these two architectural families are typically viewed as fundamentally distinct. We argue that convolution and self-attention, despite their apparent differences, can be unified within a single $k$-nearest neighbor aggregation framework. The critical insight is that both operations are special cases of neighbor selection and aggregation; convolution selects neighbors by spatial proximity, while attention selects by feature similarity, revealing they exist on a continuous spectrum. We introduce Convolutional Nearest Neighbors (ConvNN), a unified framework that formalizes this connection. Crucially, ConvNN serves as a drop-in replacement for convolutional and attention layers, enabling systematic exploration of the intermediate spectrum between these two extremes. We validate the framework’s coherence on CIFAR-10 and CIFAR-100 classification tasks across two complementary architectures: (1) Hybrid branching in VGG improves accuracy on both CIFAR datasets by combining spatial-proximity and feature-similarity selection; and (2) ConvNN in ViT outperforms standard attention and other attention variants on both datasets. Extensive ablations on $k$ values and architectural variants reveal that interpolating along this spectrum provides regularization benefits by balancing local and global receptive fields. Our work provides a unifying framework that dissolves the apparent distinction between convolution and attention, with implications for designing more principled and interpretable vision architectures.
1 Introduction
Deep learning architectures for computer vision have undergone a fundamental transformation over the past two decades. Convolutional neural networks (CNNs) dominated through pioneering architectures like AlexNet [DBLP:journals/cacm/KrizhevskySH17], VGG [DBLP:journals/corr/SimonyanZ14a], and ResNet [he2016deep]. More recently, Vision Transformers (ViT) [DBLP:conf/iclr/DosovitskiyB0WZ21] have reshaped the landscape with attention-based models achieving state-of-the-art performance across numerous vision tasks. Yet despite these successes, convolution and attention are typically treated as fundamentally distinct mechanisms suited for different situations.
Traditionally, convolution operates through local spatial aggregation, where each output is computed from a fixed spatial neighborhood. In contrast, self-attention [vaswani2017attention] enables global feature aggregation by computing similarity-based weighted combinations across all spatial positions. However, beneath these surface differences lies a common principle: both operations aggregate features from selected neighbors. The distinction is in neighbor selection–convolution selects by spatial proximity, while attention selects by learned feature similarity.
We propose Convolutional Nearest Neighbors (ConvNN), a unified framework that formalizes the interconnection between convolution and attention through $k$-nearest neighbor ($k$-NN) selection. Our key insight is that both operations are specialized instances of neighbor aggregation, differing in their selection strategy. Crucially, ConvNN can be exactly configured to recover standard convolution and exactly recover standard attention, proving both are points on a continuous spectrum.
The flexibility of ConvNN enables systematic exploration of this spectrum. Rather than choosing between convolution or attention, we can design hybrid configurations that combine spatial and feature-based neighbor selection. Our experiments on CIFAR-10 and CIFAR-100 demonstrate that these configurations achieve consistent improvements in accuracy when integrated into both CNN (VGG) and Transformer (ViT) architectures. Beyond performance gains, hybrid configurations provide interpretability by explicitly balancing local spatial features with global feature dependencies. This unified framework provides a foundation for designing next-generation operations and architectures that leverage the strengths of both mechanisms.
2 Related Works
Convolution and Attention in Vision
Convolution has dominated computer vision through architectures like AlexNet [DBLP:journals/cacm/KrizhevskySH17], VGG [DBLP:journals/corr/SimonyanZ14a], and ResNet [he2016deep], providing efficient local feature extraction with spatial inductive biases. Recently, Vision Transformers (ViT) [DBLP:conf/iclr/DosovitskiyB0WZ21] and other attention-based models [DBLP:conf/nips/ParmarRVBLS19, DBLP:conf/iccv/LiuL00W0LG21] have achieved comparable or superior performance through self-attention [vaswani2017attention], which captures long-range dependencies without spatial constraints, but at a larger computational requirement.
Several variants of the self-attention mechanism have been proposed to reduce its quadratic computational and memory complexity. Local Attention [aguilera2024local] restricts each query to attend only to nearby tokens within a fixed window, thereby exploiting locality while lowering the attention cost from $O(n^{2})$ to $O(nw)$, where $w$ is the window size. Sparse Attention [child2019generating], further generalizes this idea by enforcing structured sparsity patterns in the attention matrix, allowing each token to attend to a predefined subset of positions. These approaches preserve most of the modeling capacity of global attention while making Transformer architectures more scalable for long sequences.
Connections Between Convolution and Attention
Recent work began exploring the theoretical relationships between these seemingly distinct operations. Cordonnier et al. [DBLP:conf/iclr/CordonnierLJ20] provided evidence that self-attention with relative positional encoding can express any convolutional operation, suggesting convolution and attention may exist on a spectrum. Non-Local Neural Networks [DBLP:conf/cvpr/0004GGH18] and channel-based attention mechanisms [DBLP:journals/pami/HuSASW20] introduced global aggregation mechanisms, demonstrating that operations beyond local convolution can enhance feature learning. Furthermore, recent architectures have explicitly combined both operations. Attention Augmented Convolutional Networks[DBLP:conf/iccv/BelloZLVS19] augment convolutional layers with self-attention. CoAtNet [DBLP:conf/iccv/YuanG0ZYW21] systematically studies depthwise convolutions followed by attention blocks, showing that this combination achieves strong performance. ConViT [DBLP:conf/icml/dAscoliTLMBS21] introduces gated positional self-attention that gradually transitions from convolutional-like to attention-like behavior. Neighborhood Attention [DBLP:conf/cvpr/000100LS23] employs sliding windows to reduce attention’s computational cost while preserving spatial locality. While these hybrid architectures demonstrate the benefits of combining local and global operations, they treat convolution and attention as distinct operations combined through architectural composition rather than unified through a common theoretical principle.
$k$-Nearest Neighbors in Deep Learning
The $k$-nearest neighbor ($k$-NN) principle has recently reemerged as a valuable design component in neural network architectures. Neural Nearest Neighbors Networks [DBLP:conf/nips/Plotz018] first explored the integration of $k$-NN reasoning primarily for few-shot learning. More recently, $k$-NN mechanisms have been incorporated into Transformers to mitigate the quadratic computational and memory costs of self-attention. In particular, $k$-NN Attention for Boosting Vision Transformers (KVT) [DBLP:conf/eccv/WangWWLCLJ22] enhances Vision Transformers by introducing a $k$-nearest-neighbor attention mechanism that restricts each token’s attention to its most relevant local neighbors rather than the entire sequence. Formally, KVT introduces a $k$-NN masking function that constrains each query to attend only to its $k$ most similar keys, according to the traditional scaled dot-product similarity. The selected neighbor values are then normalized via a softmax operation, and the overall attention is computed with a reduced complexity of $O(nk\log k)$—lower than the traditional $O(n^{2}d)$ of full self-attention—where $n$ denotes the number of tokens and $d$ the embedding dimension. By adaptively selecting the $k$ most similar keys for each query in feature space, KVT effectively balances local and global context modeling, leading to better performance and faster convergence compared to standard ViTs.
Similarly, Unlimiformer [DBLP:conf/nips/Bertsch0NG23] addresses attention computation complexity by using $k$-NN indices to compute only $k$-NN distances. Routing Transformers [roy-etal-2021-efficient] rely on $k$-means clustering to model sparse attention matrices; this allows each token to attend to its $k$ closest tokens. Furthermore, in $k$-NN Attention Demystified [DBLP:conf/iclr/Haris25], a theoretical framework for $k$-NN attention is presented, reframing self-attention as expectations and leveraging lazy Gumbel sampling for efficient approximation. Memory-efficient Transformers [gupta2021memoryefficient] also addresses this complexity through a highly accurate approximation for vanilla attention with top-$k$ scores in chunks of queries. These works demonstrate the value of neighbor selection as a design principle, though they treat $k$-NN as an enhancement to existing architectures rather than a unifying framework for understanding different operations.
3 Background
3.1 Notation
Scalar variables, vectors, and matrices, are denoted in plain lowercase (e.g., $x$), bold lowercase (e.g., $\mathbf{x}$), and in plain uppercase letters (e.g. $X$), respectively. High-order tensors are denoted by bold uppercase letters (e.g., $\mathbf{X}$). The shape of a tensor is written in square brackets (e.g., the shape $\mathbf{X}\in\mathbb{R}^{a\times b\times c}$ is $[a,b,c]$). A vector of ones in $\mathbb{R}^{k}$ is denoted as $\mathbf{1}_{k}$.
We adopt Python-style indexing notation, where square brackets (e.g., $X[i,:]$) indicate row and column selection. Given a matrix $A\in\mathbb{R}^{n\times m}$, $A[i,:]$ selects the entire $i$-th row of $A$. Given a vector mask $\mathbf{m}\in\{0,1\}^{n}$, $A[\mathbf{m},:]$ selects the rows $i$ where $\mathbf{m}_{i}=0$. A feature vector (or simply feature) in a 2D tensor is a transversal tube of unitary spatial dimensions across channels.
A vector $\mathbf{x}$ is $\ell_{2}$-normalized when $|\mathbf{x}|_{2}=1$, where $|\cdot|_{2}$ denotes the $\ell_{2}$ norm. The operator $\operatorname{softmax}(\mathbf{x})$ normalizes $\mathbf{x}$ into a probability distribution, with $\operatorname{softmax}(\mathbf{x})_{i}=\exp(x_{i})/\sum_{j}\exp(x_{j})$. Finally, $\operatorname{diag}(\mathbf{v})$ denotes the diagonal matrix with the entries of $\mathbf{v}$ on its main diagonal.
3.2 $k$-Nearest Neighbors
Let $X\in\mathbb{R}^{n\times d}$ be a matrix with columns $X=\{\mathbf{x}_{i}\}_{i=1}^{n}$, where each $\mathbf{x}_{i}$ is $\ell_{2}$-normalized. In this setting the squared Euclidean distance between two points $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ is given by $2(1-\mathbf{x}_{i}^{\top}\mathbf{x}_{j})$ and their cosine similarity is $\mathbf{x}_{i}^{\top}\mathbf{x}_{j}$, which reveals that, for normalized vectors, Euclidean distance is inversely related to cosine similarity.
We define $\text{$k$-max}:\mathbb{R}^{n\times m}\to\mathbb{R}^{n\times k}$ as the operator that returns, for each row of a matrix, the $k$ largest values. Correspondingly, $\text{$k$-argmax}:\mathbb{R}^{n\times m}\to\{0,1\}^{n\times m}$ returns a multi-hot encoding [^1] A multi-hot encoded vector is a binary vector $\mathbf{v}\in\{0,1\}^{n}$ such that $\mathbf{v}_{i}=1$ if a given condition is true for index $i$, and 0 otherwise. indicating the positions of these top-$k$ values. Let $\mathcal{N}_{k}(\mathbf{x}_{i})\subset X$ be the set of $k$-nearest neighbors of $\mathbf{x}_{i}\in X$. We can compute $\mathcal{N}_{k}(\mathbf{x}_{i})$ as a submatrix of $X$ through $X[I_{i},:]\in\mathbb{R}^{k\times d}$, where indices $I$ are given by:
$I=\text{$k$-argmax}(XX^{\top})\in\mathbb{R}^{n\times n},$ (1)
and $I_{i}=I[i,:]$. In other words, $X[I_{i},:]$, selects the rows of $X$ corresponding to the $k$-nearest neighbors of $\mathbf{x}_{i}$ based its similarity to the other points.
3.3 Convolution and Attention
3.3.1 Convolution
Let $Z\in\mathbb{R}^{n\times c}$ be a 1D tensor with $c$ channels and $W\in\mathbb{R}^{k\times c}$ a convolutional kernel with bias $b\in\mathbb{R}$. A 1D convolution (Conv1D) slides $W$ across $Z$, computing local dot products and adding $b$, to produce $Y\in\mathbb{R}^{n^{\prime}\times c^{\prime}}$, where $n^{\prime}$ depends on the kernel size, stride, and padding. In 2D (Conv2D), the input $Z\in\mathbb{R}^{w\times h\times c}$ is convolved with kernels $W\in\mathbb{R}^{k\times k\times c}$ across both spatial dimensions, yielding $Y\in\mathbb{R}^{w^{\prime}\times h^{\prime}\times c^{\prime}}$. For $k=1$, the operation reduces to a pointwise linear projection across channels (a $1\times 1$ convolution).
In depthwise convolution, instead of learning kernels that span all input channels, we assign one kernel per channel. That is, each channel $Z_{i}\in\mathbb{R}^{n\times 1}$ is convolved with its own kernel $W_{i}\in\mathbb{R}^{k\times 1}$, yielding $Y_{i}\in\mathbb{R}^{n^{\prime}\times 1}$. These $c$ independent convolutions capture spatial correlations within each channel separately, without mixing information across channels.
3.3.2 Attention
Given query, key, and value matrices $Q\in\mathbb{R}^{n\times h}$, $K\in\mathbb{R}^{m\times h}$, and $V\in\mathbb{R}^{m\times v}$, respectively, the dot-product attention [vaswani2017attention] computes
$A=\operatorname{softmax}(QK^{\top})\in\mathbb{R}^{n\times m},$ (2)
and produces the output $AV\in\mathbb{R}^{n\times v}$. We omit the usual scaling factor $\sqrt{h}$ for simplicity, as it does not affect our proposed formulation in the following.
4 Convolutional Nearest Neighbors
Figure 1:
The ConvNN operation
4.1 Core Framework
We now present ConvNN, an operation that generalizes both convolution and attention, is built upon three core steps: similarity computation, neighbor selection and modulation, and weighted aggregation. Let $X\in\mathbb{R}^{n\times c}$ denote the input feature matrix. ConvNN operates as follows:
  1. 1.
    Similarity computation:
    Let $f_{Q}:\mathbb{R}^{n\times c}\to\mathbb{R}^{n\times h}$, $f_{K}:\mathbb{R}^{n\times c}\to\mathbb{R}^{n\times h}$, and be learnable linear projections. Let $Q=f_{Q}(X)$ and $K=f_{K}(X)$. After applying $\ell_{2}$ normalization to the rows of $Q$ and $K$, feature similarities are computed in $\mathbb{R}^{h}$ rather than $\mathbb{R}^{c}$, and the neighbors of each feature $\mathbf{x}_{i}$ are extracted from $V$ using these indices of these projected representations. ConvNN employs a cosine similarity between these normalized features, which means that:
    $S=QK^{\top}.$ (3)
  2. 2.
    Neighbor selection and modulation:
    Let $f_{V}:\mathbb{R}^{n\times c}\to\mathbb{R}^{n\times v}$ be a another learnable linear projection, and define $V=f_{V}(X)$. For each feature vector $\mathbf{x}_{i}\in X$, let $\mathcal{N}^{\text{proj}}_{k}(\mathbf{x}_{i})$ denote the set of $k$ rows of $V$ corresponding to the $k$ rows of $K$ most similar to $Q[i,:]$. We refer to the elements of $\mathcal{N}^{\text{proj}}_{k}(\mathbf{x}_{i})$ as the projected neighbors of $\mathbf{x}_{i}$. The similarity values and indices of these neighbors are respectively denoted as:
    $\mathbf{s}_{i}=\text{$k$-max}_{k}(S)[i,:],$ (4)
    $I_{i}=\text{$k$-argmax}_{k}(S)[i,:].$ (5)
    We introduce a placeholder function $\rho:\mathbb{R}^{k}\to\mathbb{R}^{k}$ that defines the weighting scheme, instantiated either as a constant vector ($\rho(\mathbf{z})=\mathbf{1}_{k}$) or a softmax normalization ($\rho(\mathbf{z})=\operatorname{softmax}(\mathbf{z})$). Let $S_{i}=\operatorname{diag}(\rho(\mathbf{s}_{i}))$ denote the resulting diagonal weighting matrix.
    ConvNN then constructs, for each $\mathbf{x}_{i}$, a neighborhood matrix $X_{\text{nn},i}$ defined as:
    $X_{\text{nn},i}=S_{i}\,V[I_{i},:]\in\mathbb{R}^{k\times c},$ (6)
    which represents the $k$ projected neighbors of $\mathbf{x}_{i}$, each scaled by the corresponding weight in $\rho(\mathbf{s}_{i})$ and ordered by similarity.
  1. 4.
    Weighted aggregation:
    Let $X_{\text{nn}}\in\mathbb{R}^{(kn)\times c}$ denote the concatenation of all neighborhood matrices $X_{\text{nn},i}$ along the spatial dimension. ConvNN applies a Conv1D operation (either standard or depthwise) with $c^{\prime}$ kernels, stride $k$, and kernel size $k$ to $X_{\text{nn}}$, yielding the output $Y\in\mathbb{R}^{c^{\prime}\times n}.$
For input tensors with spatial structure, such as those of shape $[c,w,h]$, ConvNN first flattens the spatial dimensions to obtain a tensor of shape $[wh,c]$, applies the procedure described above, and then reshapes the result back to $c^{\prime},w,h]$. An overview of the ConvNN is illustrated in Figure 1, and its official PyTorch [pytorch2025] code is detailed in Algorithm 1.
Algorithm 1: PyTorch code of our ConvNN module.
class ConvNN($n$.Module):
def __init__(self, k=9, rho=’softmax’):
super().__init__()
self.k, self.rho = k, rho
self.conv1d = $n$.Conv1d(kernel_size=k, stride=k)
self.w_k = $n$.Conv1d(kernel_size=1, stride=1)
self.w_q = $n$.Conv1d(kernel_size=1, stride=1)
self.w_v = $n$.Conv1d(kernel_size=1, stride=1)
def forward(self, x):
# Project the input
k, q, v = self.w_k(x), self.w_q(x), self.w_v(x)
# Compute Similarities
k_norm = F.normalize(k, p=2, dim=1)
q_norm = F.normalize(q, p=2, dim=1)
S = k_norm.T @ q_norm
# Select Neighbors
kmax, kargmax = torch.topk(S, k=self.k, dim=-1)
if self.rho == ’softmax’:
kmax = torch.softmax(kmax, dim=-1)
# Neighbor Gathering and Modulation
v = v.unsqueeze(-1).expand(*x.shape, self.k)
kmax = kmax.unsqueeze(1)
.expand(*x.shape, self.k)
kargmax = kargmax.unsqueeze(1)
.expand(*x.shape, self.k)
x_nn = kmax * torch.gather(v, dim=2,
index=kargmax)
# Neighbor Aggregation
b, c, _ = x.shape
return self.conv1d(x_nn.view(b, c, -1))
4.2 Extensions
To enhance the expressiveness of ConvNN, we introduce two extensions that enable it to interpolate between purely feature-based and spatially-aware neighbor selection.
4.2.1 Positional Encoding
To incorporate spatial structure when beneficial, ConvNN optionally concatenates positional coordinates to the input along the channel dimension. For a 1D input $[c,l]$, a single channel increasing linearly with position is added, while for a 2D input $[c,w,h]$, two channels encoding horizontal and vertical coordinates $(i,j)$ are appended. The resulting tensors have shapes $[c+1,l]$ and $[c+2,w,h]$, respectively. This positional encoding enables ConvNN to blend feature-based and spatially aware similarity computations.
4.2.2 Sparse Neighbor Search
To reduce the computational burden of $k$-NN search from $O(n^{2}\log n)$ to $O(nr\log r)$, we restrict the search to a set of $r\ll n$ candidate neighbors. We select these candidates using one of two strategies:
  • Random Selection (see Fig. 2(b)): We randomly sample $r$ features from $X$ to form $X^{\text{sparse}}$, i.e., $X^{\text{sparse}}=X[I_{r},:]$ where $I_{r}$ contains $r$ indices drawn uniformly without replacement. This strategy preserves feature diversity while maintaining a reduced search space.
  • Spatial Selection (see Fig. 2(c)): For a 1D tensor $\mathbf{X}\in\mathbb{R}^{c\times l}$ and a subsampling rate $r$, we extract a reduced tensor $\mathbf{X}^{\text{sparse}}\in\mathbb{R}^{c\times\frac{l}{r}}$ such that $\mathbf{X}^{\text{sparse}}[:,i]=\mathbf{X}[:,ri]$, effectively selecting features at regular intervals along the sequence. Analogously, for a 2D tensor $\mathbf{X}\in\mathbb{R}^{c\times w\times h}$ and $r^{\prime}=\sqrt{r}$, we extract a subtensor $\mathbf{X}^{\text{sparse}}\in\mathbb{R}^{c\times\frac{w}{r^{\prime}}\times\frac{h}{r^{\prime}}}$ such that $\mathbf{X}^{\text{sparse}}[:,i,j]=\mathbf{X}[:,r^{\prime}i,r^{\prime}j]$, thereby subsampling features at uniform spatial intervals.
(a) Input Tensor
(b) Random
(c) Spatial
Figure 2: Neighbor candidate selection strategies for $r=16$. Candidates features are highlighted in gray.
Beyond computational efficiency, this technique reduces locality bias where the closest feature vectors tend to be spatially adjacent, forcing ConvNN to learn potentially beneficial long-range relationships [DBLP:conf/iclr/KitaevKL20].
4.2.3 Hybrid Branching Layer
A key distinction between ConvNN and standard convolution lies in their receptive behavior: while convolutions aggregate information locally from spatially adjacent features, ConvNN captures long-range dependencies, allowing global context to influence each representation.
Building upon prior work that integrates local feature extraction with global feature modeling [szegedy2015going, peng2021conformer, chen2022regionvit], we introduce a hybrid branching layer that combines convolutional and ConvNN components. We first set the ConvNN projectors as identity functions. Then this layer processes the input through two parallel branches—one convolutional and one ConvNN—and fuses their outputs via a $1\times 1$ convolution. A parameter $\lambda\in[0,1]$ controls the proportion of output channels allocated to each branch, ensuring a balanced contribution between local and global processing. Both branches use the same kernel size (or neighbor count), keeping the parameter budget comparable while regulating their relative influence through $\lambda$.
5 Connection to Prior Literature
5.1 Traditional Convolution
Traditional convolution emerges as a special case of ConvNN where neighbors are selected by spatial proximity rather than feature similarity. This can be achieved by using positional encoding, choosing projections $f_{K}$ and $f_{Q}$ that filter all channels except spatial dimensions, setting $f_{V}$ as the identity function and choosing $\rho=\mathbf{1}_{k}$ (constant weighting). In this configuration, the $k$-nearest neighbors of a feature are necessarily the spatially adjacent features, reducing ConvNN to standard convolution with receptive field size given by $k$.
5.2 Attention Mechanism
Next, we show that the attention mechanism in Eq. 2 can be recovered as a special case of ConvNN by setting $k=n$, $\rho=\operatorname{softmax}$, and applying a depthwise convolution with unit weights and zero bias during the weighted aggregation phase. Consider first the multiplication $AV$ in Eq. 2. The $i$-th row of this product can be viewed as the sum of the rows of $V$, each scaled by the corresponding attention weight in $A[i,:]=\operatorname{softmax}(Q[i,:]K^{\top})$.
Under the above settings, ConvNN reproduces this computation. In Eq. 6, the modulation step scales the same rows of $V$, now arranged as $V[I^{\text{proj}}_{i},:]$, by the values in $\rho(\mathbf{s}_{i})=\operatorname{softmax}(Q[i,:]K^{\top})$. Applying a bias-free depthwise convolution with constant weights equal to one effectively performs the same summation across the scaled rows of $V$. When applied to all $i$, this process yields exactly $AV$. Furthermore, relaxing the condition $k=n$ naturally recovers the localized attention formulation proposed in KVT [DBLP:conf/eccv/WangWWLCLJ22].
6 Experimental Evaluation
Table 1: Test loss and accuracy for VGG-11 trained on CIFAR-10 and 100 when varying $\lambda$ in your hybrid branching module. Bold values indicate values that outperform those of the Conv2D baseline.
Without Positional Encoding
CIFAR-10 CIFAR-100
Layer Type Test Loss Acc (%) Test Loss Acc (%) GFLOPS
Conv2D 1.529 80.65 3.939 44.83 0.293
$\lambda$ 0.125 1.393 81.10 3.905 45.08 0.325
0.250 1.467 81.21 3.518 47.87
0.375 1.422 80.95 3.578 48.34
0.500 1.424 81.26 3.474 48.37
0.625 1.403 80.79 3.281 49.74
0.750 1.479 79.58 3.516 46.93
0.875 1.332 78.18 3.906 45.72
ConvNN 2.694 54.00% 5.011 29.50% 0.325
# Params. 130.015M 130.384M
With Positional Encoding
Layer Type CIFAR-10 CIFAR-100
Test Loss Acc (%) Test Loss Acc (%) GFLOPS
Conv2D 1.529 80.65 3.939 44.83 0.293
$\lambda$ 0.125 1.420 80.97 3.665 47.73 0.331
0.250 1.447 80.91 3.702 47.65
0.375 1.386 82.08 3.696 48.39
0.500 1.213 81.77 3.588 48.80
0.625 1.341 80.81 3.323 49.23
0.750 1.283 80.26 3.533 48.78
0.875 1.225 78.70 3.539 46.26
ConvNN 2.388 57.15 4.586 32.33 0.331
# Params. 130.015M 130.384M
We evaluate the performance of ConvNN against convolutional and attention-based variants on the CIFAR-10 and CIFAR-100 classification benchmarks [krizhevsky2009learning]. These datasets provide a controlled setting for assessing theoretical frameworks and isolating the contributions of the unified ConvNN formulation. CIFAR-10 and CIFAR-100 remain standard benchmarks for controlled architectural studies in computer vision [he2016deepresidual] and have also been recently adopted in the evaluation of transformer-based architectures [chen2023accumulated].
For convolution-based experiments, we employ a VGG convolutional neural network (VGG-11) [DBLP:journals/corr/SimonyanZ14a] architecture. For attention-based experiments, we use Vision Transformer (in its ViT-Tiny version) [DBLP:conf/iclr/DosovitskiyB0WZ21] with 12 layers, 192 hidden dimensions, 768 MLP dimensions, attention dropout and MLP dropout to 0.1. For simplicity, we decided to set the number of heads in our attention-based models to $1$. In each case, when not stated otherwise, we replace all their individual convolution or attention operations by ConvNN modules when comparing both paradigms.
We trained all models using AdamW optimizer with a learning rate of $10^{-4}$ and weight decay of 0.01 for 150 epochs with a batch size of 256. We apply gradient clipping with a maximum norm of 1.0 and use cross-entropy loss as the training criterion. We do not employ a learning rate scheduler to maintain consistency and reproducibility across experiments. All experiments are implemented in PyTorch 2.0 and executed on NVIDIA A100 PCIe 80 GB GPUs.
Figure 3: Loss comparison on VGG-11 training. In branching, we set $\lambda=0.5$. For visualization purposes, all loss curves are smoothed using a moving average with a window size of 10.
Figure 4: Impact of $k$ to ConvNN’s performance on VGG-11 training against the Conv2D baseline.
Figure 5: Impact of $r$ to ConvNN’s performance on VGG-11 training. The Random and Spatial candidate selection curves are slightly offset vertically for clarity, had the same runtime throughout the experiments.
6.1 ConvNN vs. traditional convolution
In this section, for experimental simplicity and to facilitate direct comparison with standard convolutions, we set the ConvNN projectors to identity functions and fix $\rho(\mathbf{z})=\mathbf{1}_{k}$. Only the VGG-11 architecture is evaluated in this setting.
Table 1 shows the performance of our hybrid ConvNN branching modules (see Sec. 4.2.3) for varying values of $\lambda$. In these experiments, $k=9$ is used for ConvNN, and $3\times 3$ kernels are used in the Conv2D layers. Sparse neighbor search is disabled for clarity.
Table 1 reports results when raw features are used to compute neighbor similarities, or when positional encodings are added before similarity computation (see Sec. 4.2.1). Despite the pure ConvNN configuration ($\lambda=1$) underperforming relative to the pure convolutional case ($\lambda=0$), the hybrid branching model consistently surpasses both across most of the range $0<\lambda<1$, without introducing additional learnable parameters. This effect is particularly pronounced on CIFAR-100, where the branching model improves accuracy by nearly 5%. The inclusion of positional encodings yields further, albeit modest, gains across both datasets, but is particularly noticeable in pure ConvNN.
Figure 3 presents the training and test loss trajectories for VGG-11 on CIFAR-10 and CIFAR-100 using Conv2D, ConvNN, and the branching variant with $\lambda=0.5$, where similarities are computed from raw input features. These curves highlight the advantages of the hybrid approach: while Conv2D quickly overfits, the branching configuration exhibits greater resistance to overfitting, achieves lower overall loss (around epoch 40), and converges to the best final performance by a good margin. Interestingly, even though its representational capacity is split between Conv2D and ConvNN, the branching model outperforms both, underscoring the value of jointly balancing local and global feature learning.
Figure 4 considers the branching module ($\lambda=0.5$) performances for varying a varying number of neighbors ($k$) and kernel sizes on its ConvNN and Conv2D branches, respectively. Three important observations can be made from these results: (1) the addition of a ConvNN branch consistently improves the otherwise sole Conv2D performances, (2) the choice of a higher $k$ in most cases does usually not reflect a proportionally higher performances, which means that the same accuracy increments can be achieved at a lower learnable parameter need, and (3) ConvNN’s contributions are more evident at CIFAR-100, suggesting that it may be more suitable to datasets with a lower datapoint per class ratio.
Finally, Fig. 5 presents a quantitative comparison of the sparse neighbor search strategies described in Sec. 4.2.2 against both the baseline ConvNN (which considers all features as neighbor candidates) and the standard Conv2D. As expected, the sparse variants significantly reduce ConvNN’s runtime, approaching that of Conv2D for smaller values of $r$. Moreover, increasing the number of candidate neighbors generally improves the overall performance of the proposed module, and in some cases even surpasses the configuration that uses all features. This behavior is intuitive: by enforcing neighbor selection over a spatially constrained subset, ConvNN is encouraged to capture more diverse and spatially distant relationships, thereby mitigating locality bias and complementing the locality-focused Conv2D branch within the hybrid architecture. Interestingly, the random strategy consistently outperforms both the spatial and all-feature approaches, even for small $r$. Combined with its lower computational cost, this makes it an appealing alternative to Conv2D.
Table 2: Test loss and accuracy for ViT-Tiny on CIFAR datasets. Bold values indicate best results; underline indicate second best.
Layer Type CIFAR-10 CIFAR-100
Test Loss Acc (%) GFLOPS # Params. Test Loss Acc (%) GFLOPS # Params.
Attention [vaswani2017attention] 1.558 77.41% 2.507 5.479M 3.817 50.07% 2.507 5.496M
Local-Attention [aguilera2024local] 1.626 76.51% 1.991 4.299M 3.688 49.73% 1.991 4.317M
KVT-Attention (k=9k=9[DBLP:conf/eccv/WangWWLCLJ22] 1.775 75.77% 2.508 5.479M 3.883 48.21% 2.508 5.496M
KVT-Attention (k=100k=100[DBLP:conf/eccv/WangWWLCLJ22] 1.593 76.41% 2.508 5.479M 3.738 50.34% 2.508 5.496M
Sparse-Attention [child2019generating] 2.026 72.50% 2.509 5.479M 4.565 44.62% 2.509 5.496M
ConvNN (All features) 1.978 73.67% 2.333 5.500M 3.983 48.67% 2.333 5.517M
ConvNN (Random Selection) 0.779 79.82% 2.037 5.500M 2.341 53.60% 2.037 5.517M
ConvNN (Spatial Selection) 2.737 63.01% 2.037 5.500M 5.020 38.22% 2.037 5.517M
Figure 6: Loss comparison on ViT-Tiny training. ConvNN with Random Candidate Selection sampling $r=32$. For visualization purposes, all loss curves are smoothed using a moving average with a window size of 10. Learning rate was set as $10^{-4}$.
Figure 7: Impact of $k$ to ConvNN’s performance on ViT-Tiny training against both original Attention and KVT baselines. The KVT model is also evaluated across varying $k$, and at its best performance at $k=100$.
6.2 ConvNN vs. attention
In this section, we present quantitative comparisons between the standard attention mechanism and our proposed ConvNN variant. To ensure a fair comparison, we employ the full set of linear projectors, omit positional encoding, and set $\rho=\operatorname{softmax}$. Following its connection to KVT (see Sec. 5.2), ConvNN uses a 1D depthwise convolution with weights initialized uniformly to 1 and no bias. Unless otherwise stated, we set $k=9$ and $r=32$. We further evaluate ConvNN against several representative attention variants— the aforementioned KVT Attention [DBLP:conf/eccv/WangWWLCLJ22], Local Attention [aguilera2024local], and Sparse Attention [child2019generating]—in addition to the original self-attention formulation [vaswani2017attention]. All methods are implemented and trained from scratch within the Tiny-ViT framework [DBLP:conf/iclr/DosovitskiyB0WZ21].
Table 2 compares the performance of ConvNN against the various attention mechanisms evaluated. While ConvNN underperforms when using either all features as potential neighbors or only uniformly spatially sampled candidates, it achieves substantially lower losses and faster runtimes when our random selection is employed. These gains are obtained with only a small increase in the overall parameter count. Fig. 6 further contributes to this picture as it shows how ConvNN avoids overfitting to the data, while that is observed in the traditional attention methodology in early epochs.
Figure 7 shows the impact of $k$ on the performance of our ConvNN-based attention module compared to both the baseline attention mechanism and KVT. When all features are used as potential neighbors, ConvNN achieves performance comparable to KVT. However, when employing our random neighbor selection strategy, ConvNN substantially outperforms both KVT and standard attention. As with previous observations, increasing $k$ beyond a certain value yields only marginal accuracy gains, indicating that ConvNN effectively captures the most relevant neighbors even with smaller $k$.
Figure 8 compares the random and spatial candidate selection strategies against both the baseline attention mechanism [vaswani2017attention] and the ConvNN variant using all features as potential neighbors. Across all configurations, ConvNN demonstrates faster runtimes than standard attention. With $r=16$, both sparse neighbor search strategies achieve runtimes approximately 20% faster than the baseline. In terms of classification accuracy, we observe that the spatial candidate selection performs consistently below the baseline, whereas the random selection strategy outperforms it in all scenarios—including those with the smallest candidate pools, where it also runs the fastest. These results further highlight the advantage of encouraging ConvNN to capture long-range relationships through randomized neighbor selection. This effect is less pronounced in the spatial sampling scheme, which still tends to emphasize local neighborhoods and may therefore overlook distant but informative features.
Figure 8: Impact of $r$ to ConvNN’s performance on ViT-Tiny training. The Random and Spatial candidate selection curves are slightly offset vertically for clarity, had the same runtime throughout the experiments.
7 Discussion
Our experiments reveal both the promise and limitations of unifying convolution and attention through $k$-NN. The ConvNN framework successfully demonstrates that these operations exist along a spectrum of neighbor selection strategies, with convolution representing fixed spatial selection and attention representing global, feature-based selection.
A key finding is the strong performance of random selection in ConvNN compared to considering all samples. This highlights that when employing a global receptive field, selecting the closest neighbors based on learned feature is more effective than arbitrary spatial or attention-based patterns. Notably, in the ConvNN-Attention variant, which fully utilizes its linear projectors and softmax normalization, random selection outperforms all other attention baselines, suggesting that attending to all positions may be unnecessary for learning global dependencies. This insight challenges the assumption underlying standard attention that all positions should influence the output.
Ablation studies on $k$ values reveal an important distinction between convolution and attention. While larger convolutional kernels consistently improve performance (e.g., kernel size $1\rightarrow 2\text{ or }3$), the gains from increasing $k$ in ConvNN are modest. Interestingly, in ViT-Tiny experiments, ConvNN using all features underperforms KVT, whereas the random selection variant outperforms all attention baselines. This indicates that ConvNN ’s advantage arises from flexible neighbor selection rather than the number of neighbors considered.
We also observe significant regularization effects, particularly in convolutional architectures. Adding global receptive fields through ConvNN acts as an implicit regularizer, forcing the network to balance spatial and feature-based aggregation. This prevents over-reliance on either mechanism alone, leading to more robust feature learning.
Our work also highlights several limitations. The hard selection mechanism in ConvNN limits gradient flow and may cause training instability in certain settings. Additionally, the computational cost of $k$-NN search remains higher than standard convolution, though lower than full self-attention. Extension to tasks requiring dense predictions (e.g., segmentation) will require careful consideration of how to maintain spatial correspondence during neighbor selection.
8 Conclusion
We introduced Convolutional Nearest Neighbors (ConvNN), a unified framework bridging convolution and attention through $k$-nearest neighbor selection. By showing that both operations can be viewed as forms of neighborhood aggregation—based respectively on spatial proximity and feature similarity—ConvNN offers new theoretical insight into the connection between CNNs and attention-based architectures. Experiments on CIFAR-10 and CIFAR-100 demonstrate its flexibility and strong empirical performance, achieving competitive accuracy and runtime relative to both convolutional and attention baselines. Future work will leverage the ConvNN framework to systematically analyze how different architectural components of CNNs and Transformers contribute to their representational strengths, as well as to design intermediate models that extend our branching strategy to more general convolution–attention hybrids.
9 Acknowledgments
This work was supported by Bowdoin College’s Christenfeld Summer Research Fellowship 2024, John L. Roberts Fund Fall Research Grant 2025, Last Mile Education’s Google AI 2024 award and NYC STEM Award.
We are grateful to Stephen Houser and Dj Merrill for their technical assistance with Bowdoin College’s HPC system and NVIDIA GPUs, and to the Bowdoin College Computer Science department for their sustained support throughout this project.
Code is available at the following repositories: